OpenStudy (anonymous):

A movie theater has 180 seats. If the theater sells full-price tickets at a rate of $7 per ticket, and sells exactly 70% of it's seats on each of the 3 consecutive nights, what is the theaters's total gross income from tickets for those 3 nights? $3,780 $3,024 $2,646 $1,512 $1,134

OpenStudy (anonymous):

We have 180 seats, of which 70% are full, and, for three nights we fill these at $7 each. We multiply to find the total: \[ 180\cdot.7\cdot3\cdot7=2646 \]

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

\[180*\frac{70}{100}*3*7=\text{$\$$2646} \]